# Non-linear traces of Choquet type on AF algebras

### v1

@@ 2. Choquet integrals and Choquet traces
The property (C) implies that
@@
3. Non-linear traces on finite direct sums of matrix algebras of Choquet type
@@
(C) Let $a\in A^{+}$ and $f,g\in C(\sigma(a))^{+}$ be comonotonic on $\sigma(a)$. For simplicity, we denote $\lambda_{j}(a)$ by $\lambda_{j}$ for any $j$. By comonotonicity on the finite totally ordered set $\sigma(a)$, there exists a permutation $\pi$ of $\{1,\dots,K\}$ such that
@@
and hence the map $\phi_{\alpha}$ satisfies
@@
4. Non-linear traces on AF algebras of Choquet type
@@
and put $\lambda_{K_{a}+1}(a):=0$. Since $\rho_{n,n+1}$ is a *-homomorphism, the functional calculus implies
@@
$\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b+2\varepsilon 1_{A}),$
and hence we have $\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b)$.
@@
Since $f$ and $g$ are continuous on $0,\|a\|]$, we obtain $f(a_{m})\to f(a)$ and $g(a_{m})\to g(a)$ in norm. By continuity of $\Phi_{\alpha}$, letting $m\to\infty$ yields
@@
Finally, we show that the map $\Phi_{\alpha}$ is unique. Let $\Psi:A^{+}\to[0,\infty)$ be a Choquet trace with $\Psi(p)=\alpha([p])$ for every projection $p\in A$. Fix $n$, the characterization of Choquet traces on $A_{n}$ shows that $\Psi|_{A_{n}^{+}}=\varphi_{\alpha}^{(n)}$. Hence $\Psi=f$ on $\bigcup_{n}A_{n}^{+}$, and by continuity, we conclude $\Psi=\Phi_{\alpha}$ on $A^{+}$. ∎
@@
Define $\alpha$ on classes of projections by
$\alpha([p]):=\varphi(p)\qquad(p\ \text{a projection in }A).$
@@
Let $p,q$ be projections in $A$ with $[p]=[q]$ in $\Gamma(A)$. By the fact that $\Gamma(A)$ (equivalently $K_{0}(A)^{+}$) is the inductive limit of $\Gamma(A_{n})$, there exist $n$ and projections $p^{\prime},q^{\prime}\in A_{n}$ such that
$p]=\gamma_{n}([p^{\prime}]),\qquad[q]=\gamma_{n}([q^{\prime}])\quad\text{in }\Gamma(A),$
and
$p^{\prime}]=[q^{\prime}]\quad\text{in }\Gamma(A_{n}).$
In the finite-dimensional algebra $A_{n}$, equality of classes implies that $p^{\prime}$ and $q^{\prime}$ are Murray–von Neumann equivalent, hence unitarily equivalent. Since $\varphi$ is unitarily invariant, we obtain $\varphi(p^{\prime})=\varphi(q^{\prime})$.
On the other hand, $p]=\gamma_{n}([p^{\prime}])$ means that $p$ is Murray–von Neumann equivalent to the image of $p^{\prime}$ in $A$ (and similarly for $q$ and $q^{\prime}$); in particular, $\varphi(p)=\varphi(p^{\prime})$ and $\varphi(q)=\varphi(q^{\prime})$ by unitary invariance. Therefore $\varphi(p)=\varphi(q)$, proving that $\alpha$ is well defined.

### v2

@@ 2. Choquet integrals and Choquet traces
The properties (C) and (H) imply that
@@
3. Non-linear traces of Choquet type on finite direct sums of matrix algebras
@@
(C) Let $a\in A^{+}$ and $f,g\in C(\sigma(a))^{+}$ be comonotonic on $\sigma(a)$. For simplicity, we denote $\lambda_{j}(a)$ by $\lambda_{j}$ for any $j$. By comonotonicity on the finite set $\sigma(a)$, there exists a permutation $\pi$ of $\{1,\dots,K\}$ such that
@@
and hence the map $\varphi_{\alpha}$ satisfies
@@
4. Non-linear traces of Choquet type on AF algebras
@@
and put $\lambda_{K_{a}+1}(a):=0$. Since $\rho_{n,n+1}$ is a unital *-homomorphism, we have
@@
$\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b+2\varepsilon 1_{A}).$
Since $\varepsilon$ is arbitrary, we have $\Phi_{\alpha}(a)\leq\Phi_{\alpha}(b)$.
@@
Since $f$ and $g$ are continuous on $\sigma(a)$, we obtain $f(a_{m})\to f(a)$ and $g(a_{m})\to g(a)$ in norm. By continuity of $\Phi_{\alpha}$, letting $m\to\infty$ yields
@@
Finally, we show that the map $\Phi_{\alpha}$ is unique. Let $\Psi:A^{+}\to[0,\infty)$ be a Choquet trace with $\Psi(p)=\alpha([p])$ for every projection $p\in A$. By Lemma 2.2, $\Psi$ is Lipschitz continuous. Fix $n$, the characterization of Choquet traces on $A_{n}$ shows that $\Psi|_{A_{n}^{+}}=\varphi_{\alpha}^{(n)}$. Hence $\Psi=f$ on $\bigcup_{n}A_{n}^{+}$, and by continuity, we conclude $\Psi=\Phi_{\alpha}$ on $A^{+}$. ∎
@@
Define $\alpha$ on $\Gamma(A)$ by
$\alpha([p]):=\varphi(p)\qquad\text{for every }[p]\in\Gamma(A).$
@@
Let $p,q\in A$ be Murray–von Neumann equivalent projections. Since $A$ is an AF algebra, we have
$1-p]=[1-q]\quad\text{in }K_{0}(A).$
Hence $p$ and $q$ are unitarily equivalent. Therefore, the map $\alpha$ is well-defined.